Solve each of the following equations for x, and check your results
2(x -3) - 3(x + 5) = 3(x - 2) – 7 = x = -2

Answer by funmath(2933)

2(x -3) - 3(x + 5) = 3(x - 2) – 7 = x = -2
This is not typed in correctly. The way you have it now, can't be solved. Send me the problems separated by a simicolom ; and I'll show you how to solve them.
Your goal is to get the x on one side of the = sign all by itself within the rules of algebra. Whatever you do to one side of the equal sign, you also do to the other.
You check your answers by substituting your answer for x in for x in the original equation, simplifying the opposing sides and seeing if they equal.
Happy Calculating!!!